OpenStudy (zyberg):

Find the last three digits of 9^8^7^6^5^4^9.

OpenStudy (zyberg):

So, I need to find mod 125 and 8. For 8 it's easy - it's 1. However, what is for 125? I don't understand how to solve it from: 6^5^4^9 mod 16. (used totient function to climb the tower of exponents)

ganeshie8 (ganeshie8):

Is this from textbook or you have cooked it up ?

OpenStudy (zyberg):

This one is textbook one. ;)

OpenStudy (zyberg):

Hm, perhaps I noticed something? 6 is same as -10 in mod 16, and 5^4^9 is an odd number, so -10^5^4^9 is congruent to -10 or to 6 in mod 16? Is it right?

OpenStudy (zyberg):

After following that and doing some ugly things I arrived at an answer of 66, is it right?

OpenStudy (zyberg):

And last three digits are 441?

ganeshie8 (ganeshie8):

Can you show me how you climbed the tower ? I don't get why you're considering mod 16

OpenStudy (zyberg):

Well, we need 9^8^7^6^5^4^9 mod 125. phi(125)= 100, so 8^7^6^5^4^9 mod 100; phi(100)=40, so 7^6^5^4^9 mod 40; phi(40)=16, so 6^5^4^9 mod 16? Or is it wrong?

OpenStudy (zyberg):

Ohh I see where I made a mistake.

OpenStudy (zyberg):

That's sad, I thought that everything was that easy :( (8 and 100 aren't coprime)

ganeshie8 (ganeshie8):

Very good attempt though Let's see how else can we work this

OpenStudy (zyberg):

But hm, we could divide that mod 100 into mod 25 and mod 4. So, mod 4 of 8^7^6^5^4^9 is 0 and 25 and 8 are coprime.

ganeshie8 (ganeshie8):

Definitely, looks lengthy hmm

OpenStudy (zyberg):

and 7 and 20 are coprime, so climbing even further.

ganeshie8 (ganeshie8):

I'll get back on this after dinner. 1 hour or so

OpenStudy (zyberg):

Okay!

OpenStudy (zyberg):

Found that the answer is 521, going to solve it once more with more attention to see if it's really 521.

OpenStudy (kainui):

Here's sorta some ideas I cooked up. First this is what we want to compute: \[9^{8^{7^{6^{5^{4^9}}}}} \mod 1000\] So I thought, well if there's some value n that isn't too hard to find that this is true: \[9^n \equiv 9 \mod 1000\] Then that means we only have to find this value: \[8^{7^{6^{5^{4^9}}}} \mod n\] Now it seems we can keep repeating the process as far deep in as we need, since we can correspondingly look at \[8^k \equiv 8 \mod n\] and \[7^{6^{5^{4^9}}} \mod k\] etc to keep going deeper if we like. I dunno if this is useful or not just an interesting observation.

OpenStudy (zyberg):

Well, yes. That's what I have been doing with Euler phi function there. It certainly is one way to solve this, not sure if the most efficient one tho ;) (took me around 15 minutes of constant multiplying and dividing numbers)

OpenStudy (zyberg):

Finished checking my work, it's 721, not 521. Ehh. Certainly took a lot of work. :/

OpenStudy (kainui):

Oh beats me, I didn't read anything, I don't even remember how to use the phi function to solve these kinds of things anyways haha.

ganeshie8 (ganeshie8):

Hey still here ?

ganeshie8 (ganeshie8):

You may use your work from your previous problem here : 7^6^5^4^9 = 1000k+601

OpenStudy (zyberg):

Thank you! Already solved it in a long way, and tried to solve it in the last problems way. :)

ganeshie8 (ganeshie8):

Good